OpenStudy (anonymous):

Hello guys I need help with this equation Simplify: 8 -2/3(the -2/3 is an exponent)

terenzreignz (terenzreignz):

LOL sure :) First deal with the 'negativeness' of the exponent, that brings it to the denominator (with a positive exponent) \[\Large a^{-m}= \frac1{a^m}\]

terenzreignz (terenzreignz):

\[\huge \frac{1}{8^{\frac23}}\] Catch me so far?

OpenStudy (anonymous):

yes I got that part but I flipped it so the 1 would be at the bottom after that I am lost lol

terenzreignz (terenzreignz):

Okay, so, you also should note that... \[\huge a^{\frac{m}n}= \sqrt[n]{a^m}\]

terenzreignz (terenzreignz):

= \[\huge [\sqrt[n]a]^m\]

terenzreignz (terenzreignz):

So that makes it... \[\huge = \frac1{[\color{red}{\sqrt[3]8}]^2}\]

OpenStudy (anonymous):

Ok I am confused with = [a − − √ n ] m why did you do that?

terenzreignz (terenzreignz):

No, that's kinda the definition of a fractional exponent... \[\huge a^{\frac{m}n}= \sqrt[n]{a^m}=[\sqrt[n]a]^m\] It becomes a radical, with the index being the denominator, and the entire thing is raised to the numerator...

terenzreignz (terenzreignz):

<ehem> *index being the denominator* \[\huge a^{\frac{m}{\color{red}n}}= \sqrt[\color{red}n]{a^m}=[\sqrt[\color{red}n]a]^m\] *and the entire thing is raised to the numerator* \[\huge a^{\frac{\color{blue}m}n}= \sqrt[n]{a^{\color{blue}m}}=[\sqrt[n]a]^{\color{blue}m}\]

OpenStudy (anonymous):

so it would be (3/8)-2 (I have no idea how to add the brackets im sorry)

terenzreignz (terenzreignz):

just stick to this:\[\huge = \frac1{[\color{red}{\sqrt[3]8}]^2}\]

OpenStudy (anonymous):

also there is an answer key to this which is a) 1/4 b)4 c) 2 |/3 d ) 16 e) -1/4

terenzreignz (terenzreignz):

no need for that, what is the cube root of 8?

OpenStudy (anonymous):

c has that weird dividion sigh I forgot the name for it

terenzreignz (terenzreignz):

square root sign or radical sign

OpenStudy (anonymous):

64

OpenStudy (anonymous):

I think it is a radical

terenzreignz (terenzreignz):

I said cube root of 8 :P

OpenStudy (anonymous):

ohh 512

terenzreignz (terenzreignz):

Still no... I said cube ROOT not cube. Means which number, when cubed, gives you 8?

OpenStudy (anonymous):

2

terenzreignz (terenzreignz):

Good :) So...\[\large \sqrt[3]8 = 2\] Consequently, our expression becomes... \[\huge = \frac1{[\color{red}{2}]^2}\] Care to finish it off? :P

OpenStudy (anonymous):

wait what about the negative 2?

terenzreignz (terenzreignz):

There is no negative 2?

OpenStudy (anonymous):

yes there is the problem is 8 -2/3

terenzreignz (terenzreignz):

We already got rid of that negative by bringing the 8 to the denominator remember? Look at the first thing I said ^^

terenzreignz (terenzreignz):

And you said you got that part -.-

OpenStudy (anonymous):

lol I just need to try to remember that's all and how come the 2- becomes a positive if I bring it down?

OpenStudy (anonymous):

-2*

terenzreignz (terenzreignz):

Yes, it does. Exponents change sign when you switch them from numerators and denominators...

OpenStudy (anonymous):

and numerators and denominators are? (im sorry im a bit rusty when it comes to math)

terenzreignz (terenzreignz):

\[\huge \frac{numerator}{denominator}\]

OpenStudy (anonymous):

so if the problem was 8 2/3 would I have to change it to a negative when I switch them?

terenzreignz (terenzreignz):

Yes, if ever you want to bring that 8^(2/3) to the denominator, the exponent becomes negative, though I see no reason to do that, we generally switch from numerator to denominator to get rid of negative exponents and not the other way around...

OpenStudy (anonymous):

Well okay thank you very much for the help terenzreignz :o)

terenzreignz (terenzreignz):

no problem :)

terenzreignz (terenzreignz):

And it's Terence btw :P

OpenStudy (anonymous):

Well thank you Terence and my name is Kay :0)

terenzreignz (terenzreignz):

Noted :)